# Weight Ensembling Improves Reasoning in Language Models

### v3

@@ 8 Acknowledgements
We’d like to thank Aviral Kumar, Sean Welleck, Amrith Setlur and Yiding Jiang for insightful discussions about test-time scaling and reinforcement learning. We’d also like to thank Alex Li, Sachin Goyal, and Jacob Springer for their meaningful contribution to our figures and literature review. We gratefully acknowledge support from Apple, Google, Cisco, OpenAI, NSF, Okawa foundation, the AI2050 program at Schmidt Sciences (Grant ##\##G2264481), and Bosch Center for AI.
@@ C.2 Notations and Setup
Formally we consider the following setting. We consider a $K+1$-armed bandit, with arms $\{1,2,\dots,K+1\}$. Arms $1,\dots,K$ are “good,” each yielding reward $1$, and the other arm is “bad,” yielding reward 00. We use a softmax parameterization:
@@ C.3 Implicit Diversity Collapse without KL regularization
$\lim_{t\to\infty}\max_{i\in[K]}p_{i}^{(t)}=1.$
Thus, the policy asymptotically collapses to a single-arm strategy.
@@
We define two properties of policy update rule that will lead to diversity collapse
@@
  2. 2.
    A policy update rule is said to be self-enforcing stochastic if it is self-enforcing and there exists constants $C_{1},C_{2}>0$ such that for any $\epsilon>0$, whenever the current policy satisfies $\max_{i\in[K]}p_{i}^{(t)}\in[1/2K,1-\epsilon]$ (i.e., no single good arm dominates), for $i^{*}=\arg\max_{i\in[K]}p_{i}^{(t)}$ the conditional second moment of the parameter updates for every good arm $i\in[K+1]$ and $i\neq i^{*}$satisfies:
@@
Using Lemma C.8 and C.9, we can show that REINFORCE and GRPO are self-enforcing stochastic policy update rules when bad arm probability is lower than $1/2$. The proof is then complete. ∎
@@
Under the GRPO algorithm without KL regularization ($\beta=0$),$\lim_{t\to\infty}p_{K+1}^{(t)}=0$ almost surely.
@@
Consider a policy update process that is self-enforcing stochastic (Definition C.2), then $\lim_{t\to\infty}\max_{i\in[K]}p_{i}^{(t)}=1$ almost surely.
@@
Assume that $i^{*}=\arg\max_{i\in[K]}p_{i}^{(t)}$ and $\epsilon=1-p_{i^{*}}^{(t)}$. We will reuse the notation $C$ in Definitio C.2.
Lemma C.6.
@@
For any $i\neq i^{*}$ and $i\in[K+1]$, there exists constant $B$ depend on $C_{1},C_{2}$, such that $\mathbb{E}[\exp(\theta_{i}^{t+1}-\theta_{i^{*}}^{t+1})\mid\theta^{(t)}]<\exp(\theta_{i}^{t}-\theta_{i^{*}}^{t})/(1+B\epsilon^{2})$
Proof.
Define $u_{i}=(\theta_{i^{*}}^{t+1}-\theta_{i^{*}}^{t})-(\theta_{i}^{t+1}-\theta_{i}^{t})$
We know that
$|u_{i}|<2C_{2}.$
Therefore according to Lemma C.11, there exists constant $A_{2C_{2}}>0$, such that
$\exp(u_{i})\geq 1+u_{i}+A_{2C_{2}}u_{i}^{2}.$
This then implies
$\mathbb{E}[\exp(u_{i})]\geq 1+\mathbb{E}[u_{i}]+A_{2C_{2}}\mathbb{E}[u_{i}^{2}].$
Here $\mathbb{E}[u_{i}]>0$ for $i\in[K]$ because the policy update ruleis self-enforcing and $\mathbb{E}[u_{i}]>0$ for $i=K+1$ because
$\mathbb{E}[\theta_{i^{*}}^{(t+1)}-\theta_{i^{*}}^{(t)}]>0$
$\mathbb{E}[\theta_{K+1}^{(t+1)}-\theta_{K+1}^{(t)}]<0$
Further,
$\mathbb{E}[u_{i}^{2}]\geq C_{1}\epsilon^{2}$
because the update rule is self-enforcing stochastic.
Therefore,
$\mathbb{E}[\exp(u_{i})]\geq 1+A_{2C_{2}}C_{1}\epsilon^{2}.$
Lemma C.7.
For the same constant $B$ defined in Lemma C.6,
$\mathbb{E}[\frac{1-p_{i^{*}}^{(t+1)}}{p_{i^{*}}^{(t+1)}}\mid\theta^{(t)}]<\frac{1-p_{i^{*}}^{(t)}}{p_{i^{*}}^{(t)}}/(1+B\epsilon^{2})$
Proof.
Summing the inequality for all $i\neq i^{*}$ in Lemma C.6 gives this exact inequality. ∎
Consider $p^{(t)}=\max_{i\in[K]}p_{i}^{(t)}$, according to Lemma C.7,
$\mathbb{E}[\frac{1-p^{(t+1)}}{p^{(t+1)}}\mid\theta^{(t)}]<\mathbb{E}[\frac{1-p_{i^{*}}^{(t+1)}}{p_{i^{*}}^{(t+1)}}\mid\theta^{(t)}]<\frac{1-p^{(t)}}{p^{(t)}}/(1+B\epsilon^{2}).$
Define $u_{t}=\frac{1-p^{(t)}}{p^{(t)}}>0$, then $p^{(t)}=\frac{1}{1+u_{t}}$, we can the rewrite the above equation as
$E[u_{t+1}|\theta_{t}]\leq u_{t}\frac{(1+u_{t})^{2}}{(1+u_{t})^{2}+Bu_{t}^{2}}\leq u_{t}.$
This then suggests $u_{t}$ is a positive super-martingale, according to Doob’s first convergence theorem, $u_{t}$ converges to a non-negative random variable $u_{\infty}$ almost surely. We can then show that
$u_{\infty}\leq u_{\infty}\frac{(1+u_{\infty})^{2}}{(1+u_{\infty})^{2}+Bu_{\infty}^{2}}.$
There fore $u_{\infty}=0$. Therefore, $p^{(t)}$ converges to 1 almost surely. ∎
@@ Appendix D Open-Thoughts Evaluation
We full finetune Qwen2.5-7B-Instruct over OpenThoughts-114k for 5 epochs using BF16 and AdamW with hyperparameters learning-rate=1e-5, batch-size=128, warmup=150 steps. We sample 40 reasoning traces with temperature set to $0.7$ for each of the 30 problems in AIME24. Then we evaluate the following quantities.
Figure 10: Pass@K Evaluated on AIME24 over OpenThoughts-114K SFT checkpoints. We plot the expected $\mathrm{Pass@K}$$\pm$ SD. Note that the rate at which Pass@K slows down while $\mathrm{Pass@1}$ improves at a constant rate. Furthermore, note that the confidence interval of $\mathrm{Pass@1}$ widens, meaning the variance increases during SFT.
@@
Figure 12: We plot the average number of unique answers sampled over the total number samples i.e. |$\frac{\left|\{y_{i}\}_{i=1}^{n}\right|}{n}$. Model samples less diverse number of answers as SFT progresses.
@@ Appendix F Measuring Diversity of Traces
The cardinality or number of unique answers in the set of all model outputs |$\left|\{\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{n}\}\right|$ over the total number of traces.

### v4

@@ 8  Acknowledgements
We’d like to thank Aviral Kumar, Sean Welleck, Amrith Setlur and Yiding Jiang for insightful discussions about test-time scaling and reinforcement learning. We’d also like to thank Alex Li, Sachin Goyal, and Jacob Springer for their meaningful contribution to our figures and literature review. We gratefully acknowledge support from Apple, Google, Cisco, OpenAI, NSF, Okawa foundation, the AI2050 program at Schmidt Sciences (Grant #\#G2264481), and Bosch Center for AI.
@@ C.2 Notations and Setup
Formally we consider the following setting. We consider a $K+1$-armed bandit, with arms $\{1,2,\dots,K+1\}$. Arms $1,\dots,K$ are “good,” each yielding reward $1$, and the other arm is “bad,” yielding reward 0. We use a softmax parameterization:
@@ C.3 Implicit Diversity Collapse without KL regularization
$\limsup_{t\to\infty}\max_{i\in[K]}p_{i}^{(t)}=1.$
Thus, the policy collapses to a single-arm strategy during training.
@@
We define three properties of policy update rule that will lead to diversity collapse
@@
    The policy update takes the form of $\sum_{k=1}^{B}A_{k}\nabla\log p_{i}(\theta_{i_{k}})$ where $i_{k}$ is the $k$-th sampled arm in the batch and $A_{k}$ is a function determined by (i) the sum of reward $\sum_{i=1}^{K}r_{i_{k}}$ with in the batch ; (ii) the reward $r_{i_{k}}$ and (iii) the batch size $B$.
  2. 2.
@@
  3. 3.
    A policy update rule is said to be self-enforcing stochastic if it is self-enforcing and there exists constants $C_{1},C_{2}>0$ such that for any $\epsilon>0$, whenever the current policy satisfies $\max_{i\in[K]}p_{i}^{(t)}\in[1/2K,1-\epsilon]$ (i.e., no single good arm dominates), for $i^{*}=\arg\max_{i\in[K]}p_{i}^{(t)}$ the conditional second moment of the parameter updates for every arm $i\in[K+1]$ and $i\neq i^{*}$satisfies:
@@
Using Lemma C.6 and C.7, we can show that REINFORCE and GRPO are self-enforcing stochastic policy update rules when bad arm probability is lower than $1/2$. The proof is then complete. ∎
@@
[…]
@@
Consider a policy update process that is self-enforcing stochastic (Definition C.2), then $\limsup_{t\to\infty}\max_{i\in[K]}p_{i}^{(t)}=1$ almost surely.
@@
We will inductively prove that for different $K$ the following induction hypotheses, for any $\epsilon,\delta>0$, there exists $T_{\epsilon,\delta,K}>0$,
$\Pr(\max_{t<T_{\epsilon,\delta,K}}\max_{i\in[K]}p_{i}^{(t)}<1-\epsilon)<\delta.$
We first consider the case where $K=2$.
Consider the stopping time,
$\tau_{\epsilon}=\arg\min_{t}\max_{i\in[K]}p_{i}^{(t)}>1-\epsilon$
For any $\mathcal{I}=\{1,2\}$,define $\Delta_{\mathcal{I}}^{t}=\max_{j\in[K]}\theta_{j}^{t}-\min_{j\in\mathcal{I}}\theta_{i}^{t}$.
Assume $\theta_{i*}^{t}=\max_{j\in[K]}\theta_{j}^{t}$, because $|\mathcal{I}|\geq 2$, there exists $i\neq i^{*}$, $\min_{j\in\mathcal{I}}\theta_{i}^{t}>0$. We will show three properties of $\Delta_{\mathcal{I}}^{t}$
First $\Delta_{\mathcal{I}}^{(t)}$ is a submartingale defined on the filtration of the distribution of $\theta^{(t)}$ because
$\mathbb{E}[\Delta_{\mathcal{I}}^{(t)}\mid\theta_{t}]-\Delta_{\mathcal{I}}^{(t-1)}>\mathbb{E}[(\theta_{i^{*}}^{t+1}-\theta_{i^{*}}^{t})-(\theta_{i}^{t+1}-\theta_{i}^{t})\mid\theta_{t}]>0.$
as the policy is self-enforcing.
[…]
$|\max_{j\in[K]}\theta_{j}^{t+1}-\max_{j\in[K]}\theta_{j}^{t}|<C_{2}.$
$|\min_{j\in\mathcal{I}}\theta_{j}^{t+1}-\max_{j\in\mathcal{I}}\theta_{j}^{t}|<C_{2}.$
Furthermore, the second-momentum of $\Delta_{\mathcal{I}}^{(t)}$ needs to increase with $t$ by a constant for any $t<\tau_{\epsilon}$.
$\mathbb{E}[(\Delta_{\mathcal{I}}^{(t+1)})^{2}\mid\theta_{t}]$ $\geq(\Delta_{\mathcal{I}}^{(t)})^{2}+\mathbb{E}[(\Delta_{\mathcal{I}}^{(t+1)}-\Delta_{\mathcal{I}}^{(t)}))^{2}\mid\theta_{t}]$
$\geq(\Delta_{\mathcal{I}}^{(t)})^{2}+C_{1}\epsilon^{2}.$
When $t<\tau_{\epsilon}$, it holds that $\Delta_{\mathcal{I}}^{(t)}<\log\frac{2}{\epsilon}$, otherwise we can prove that
$\max_{i,j\in\{1,2\}}p_{i}/p_{j}=\exp(\Delta_{\mathcal{I}}^{(t)})>\frac{2-2\epsilon}{\epsilon}.\implies\max_{i\in\{1,2\}}p_{i}>1-\epsilon.$
This is a contradiction. Further, by Martingale inequality, we have that
$\mathbb{E}[\left(\Delta^{\min\{t,\tau_{\epsilon}\}}\right)^{2}]>\mathbb{E}[\left(\Delta^{0}\right)^{2}]+C_{1}\epsilon^{2}\mathbb{E}[\min\{t,\tau_{\epsilon}\}]$
Further, as $\Delta^{t}$ has bounded growth, we have that
$\mathbb{E}[\left(\Delta^{\min\{t,\tau_{\epsilon}\}}\right)^{2}]<(\log\frac{2}{\epsilon}+2C_{2})^{2}.$
This implies $\mathbb{E}[\min\{t,\tau_{\epsilon}\}]<\frac{(\log\frac{2}{\epsilon}+2C_{2})^{2}}{C_{1}\epsilon^{2}}$ for all $t$, this implies
$\mathbb{E}[{\tau_{\epsilon}}]<\frac{(\log\frac{2}{\epsilon}+2C_{2})^{2}}{C_{1}\epsilon^{2}}.$
Further, by Markov inequality, if we choose
$T_{\epsilon,\delta,2}=\frac{(\log\frac{2}{\epsilon}+2C_{2})^{2}}{C_{1}\epsilon^{2}\delta}.$
then,
$\Pr(\tau_{\epsilon}>T_{\epsilon,\delta,2})<\frac{\mathbb{E}[{\tau_{\epsilon}}]}{T_{\epsilon,\delta,2}}<\delta.$
This concludes the proof for $K=2$.
Now assuming the result holds for K1K-1 and consider the case for $K$, First, we choose a small enough constant $C_{\delta,\epsilon,K,N}>0$, such that when $p_{K-1}^{(0)}<C_{\delta,\epsilon,K,N}$, the following two random processes are close:
  • Running the algorithm for $N$ steps on the $K$ arms bandit yields $\theta_{i}^{(t)},i\in[K]$
  • Running the algorithm for $N$ steps on a K1K-1 arms bandit yields $\tilde{\theta}_{i}^{(t)},i\in[K-1]$ with $\tilde{\theta}_{i}^{(0)}=\theta_{i}^{(0)},i<K-1$ and $\tilde{\theta}_{K-1}^{(0)}=\theta_{K}(0)$
@@
and there exists a joint measure on $\theta$ and $\tilde{\theta}$ such that
$\forall i\in[K-2],t<N,\Pr(|p_{i}^{t}-\tilde{p}_{i}^{t}|>\epsilon/2)<\delta/6.$
$\Pr(|p_{K}^{t}-\tilde{p}_{K-1}^{t}|>\epsilon/2)<\delta/6.$
$\Pr(|p_{K}^{t}-p_{K}^{0}|>\epsilon/2)<\delta/6.$
This joint measure is constructed by choosing the corresponding arm for two process at each sampling step as long as the sampled arm is not $K$ and uses the uniform convergence on $\nabla\log_{\theta}p_{i}$. Now following the same argument at $K=2$, we can show that there exists $\tilde{T}_{\epsilon,\delta,K}$ such that
$\Pr(\exists t<\tilde{T}_{\epsilon,\delta,K},\min_{t\in[K]}p_{t}<C_{\delta,\epsilon,K,T_{\epsilon/2,\delta/2,K-1}})>1-\delta/2.$
Then we can invoke the induction hypothesis and uses the coupling shown above to show that if we choose $T_{\epsilon,\delta,K}=\tilde{T}_{\epsilon,\delta,K}+T_{\epsilon/2,\delta/2,K-1}$, then there exists a time step that one arm has probability higher than $1-\epsilon$ with probability at least $1-\delta$.
@@ Appendix D Open-Thoughts Evaluation
We finetune Qwen2.5-7B-Instruct over OpenThoughts-114k for 5 epochs using BF16 and AdamW and hyperparameters lr=1e-5, bs=128, warmup=150 steps. We sample 40 reasoning traces with temperature set to $0.7$ for each of the 30 problems in AIME24. Then we evaluate the following quantities.
Figure 10: Pass@K Evaluated on AIME24 over OpenThoughts-114K SFT checkpoints. We plot the expected $\mathrm{Pass@K}$ $\pm$ SD. Note that improvements in Pass@K slows down while $\mathrm{Pass@1}$ improves at a constant rate. Furthermore, the confidence interval of $\mathrm{Pass@1}$ widens, meaning the variance increases during SFT.
@@
Figure 12: We plot the average number of unique answers sampled over the total number samples i.e. $\left|\{y_{i}\}_{i=1}^{n}\right|/{n}$. Model samples less diverse number of answers as SFT progresses.
@@ Appendix F Measuring Diversity of Traces
The cardinality or number of unique answers in the set of all model outputs $\left|\{\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{n}\}\right|$ over the total number of traces.
@@ F.2 How well do token-level diverse decoding strategies compare with optimal strategy with oracle?
F.3 Diversity Comparison Between SFT and WiSE-FT
Figure 21: Operation, Semantic, and Answer Diversity of Gemma-2-2B checkpoints of SFT over GSM8K versus the corresponding WiSE-FT variants (with the earliest checkpoint). We decode with temperature set to 1.0.
Figure 22: Operation, Semantic, and Answer Diversity of Gemma-2-2B checkpoints of SFT over GSM8K versus the corresponding WiSE-FT variants (with the earliest checkpoint). We decode with temperature set to 1.6.
@@ Appendix G Best of K Evaluation
Appendix H Diversity Collapse and WiSE-FT Results for the Coding Task
To test whether coding tasks exhibit the same diversity collapse observed in reasoning benchmarks, we fine-tuned the Qwen2.5-coder-0.5B model for 10 epochs on the Magicoder-Evol-Instruct-110K dataset, following the Stage 2 SFT recipe from OpenCoder LLM. We then applied WiSE-FT by interpolating the weights of the second SFT checkpoint with the initial model using interpolation ratio 0.5. Both the original SFT checkpoints and their WiSE-FT counterparts were evaluated on HumanEval for pass@k.
We found that, much like in mathematical reasoning tasks, SFT on coding data indeed suffers from diversity collapse: although pass@1 steadily improves over epochs, pass@k begins to deteriorate. And WiSE-FT still improves performance and mitigates the diversity collapse.
Figure 26: Pass@K performance of SFT checkpoints on HumanEval (temperature = 1.0).
Figure 27: Comparison of pass@K for SFT checkpoints and their WiSE-FT counterparts at k = 1, 16, 64.
Figure 28: Pass@K performance of the final SFT checkpoint versus its WiSE-FT variant.